# April 4, 2025

### v2

@@ 4. Bounding the Probabilities of Causation
$\mathrm{PS}$ =$=\mathbb{P}\{Y(1)=1\mid Y=0,D=0\},$
$\mathrm{PN}$ =$=\mathbb{P}\{Y(0)=0\mid Y=1,D=1\},$
$\mathrm{PNS}$ =$=\mathbb{P}\{Y(1)=1,Y(0)=0\},$
@@
$\mathrm{PS}$ =$=\frac{\mathbb{P}\{\mathrm{TP},D=0\}}{\mathbb{P}\{\mathrm{TP},D=0\}+\mathbb{P}\{\mathrm{NP},D=0\}},$
$\mathrm{PN}$ =$=\frac{\mathbb{P}\{\mathrm{TP},D=1\}}{\mathbb{P}\{\mathrm{TP},D=1\}+\mathbb{P}\{\mathrm{AP},D=1\}},$
$\mathrm{PNS}$ =$=\mathbb{P}\{\mathrm{TP}\}.$
Under assumptions A, C and B, we have that $\theta=\mathrm{PS}$, $\theta^{(r)}=\mathrm{PN}$, and $\mathrm{PNS}=\mathbb{P}(Y=1\mid D=1)-\mathbb{P}(Y=1\mid D=0)$, as initially noted by Pearl (1999). More interestingly, under assumptions A, B and D, we have the following results.
@@
Suppose that assumptions A, B and D hold. Then, the sharp identified intervals on $\mathrm{PS}$, $\mathrm{PN}$, and $\mathrm{PNS}$ are given by
$0\leq\mathrm{PS}\leq\theta_{U},\;0\leq\mathrm{PN}\leq\theta^{(r)}_{U},\;\text{ and }\;0\leq\mathrm{PNS}\leq\mathbb{P}(Y=1\mid D=1)-\mathbb{P}(Y=1\mid D=0).$
@@
Theorem 2 therefore shows that under assumptions A, B and D, the bounds on $\mathrm{PS}$ and $\mathrm{PN}$ coincide with those on $\theta$ and $\theta^{(r)}$, respectively. In other words, $\theta_{U}$ and $\theta^{(r)}_{U}$ also serve as the sharp upper bounds on $\mathrm{PS}$ and $\mathrm{PN}$, respectively. In addition, it confirms that $\mathrm{PNS}$ shares the same bounds as $\mathbb{P}\{\mathrm{TP}\}$, since both correspond to the average treatment effect.
@@ 5. Empirical Example
In this section, we revisit LP’s findings to reevaluate how effectively TV debates persuade (undecided) voters. The debates examined in LP took place between 5 and 44 days (on average 24 days) before the election, precisely when vote choice consistency tends to grow most swiftly. By employing an event-study design and capitalizing on the fact that debates occurred at varying times across different elections, LP control for the number of days preceding each election using daily fixed effects to distinguish the impact of the debates from general time trends. Interestingly, LP find no significant differences in how voters’ choices evolved before and after debate periods, rejecting any effect larger than $0.5$ percentage points. This null result holds for diverse voter types, different election contexts, and even tight, uncertain races.
In LP’s study, each observation corresponds to a respondent $\times$ debate $\times$ election, amounting to 331,000 total observations. The main specification of the regression model used in LP is (in their notation):
$Y_{it}^{d}=\sum_{k=-3}^{-1}\mu_{k}+\sum_{k=1}^{3}\mu_{k}+\mu_{4-}+\mu_{4+}+\sum_{t=-1}^{-60}\beta_{t}D_{t}+\alpha^{d}+W_{it}^{\prime}\lambda+u_{it}^{d},$ (2)
where $Y_{it}^{d}$ is the outcome for respondent $i$, surveyed $t$ days before the election linked to debate $d$. $D_{t}$’s, not to be confused with the treatment variable in our notation, are indicators to control for the number of days prior to the election, and the vector $W_{it}$ represents covariates such as day-of-the-week, postelectoral survey lag fixed effects, and sociodemographic characteristics. The terms $\mu_{k}(-3\leq k\leq 3)$ represent (the effects of) indicator variables for the number of days relative to the debate; $\mu_{4^{-}}$ and $\mu_{4_{+}}$ are indicators for being surveyed four or more days before or after the debate, respectively; and $\alpha^{d}$ are debate $\times$ election fixed effects.
The key coefficients of interest are $\mu_{1},\mu_{2}$, and $\mu_{3}$, which capture the effect of the debate one to three days afterward, relative to the omitted category $\mu_{0}$. They use the debate day as the reference because debates happen in the evening. They assume that any pre-trend appearing earlier than the fourth day preceding a debate or any effect emerging after the fourth day following a debate is adequately captured by the fixed effects $\mu_{4^{-}}$ and $\mu_{4^{+}}$.
Their identifying assumption is the one for the validity of the linear regression model in equation 2, which essentially requires that the date on which a debate occurred be unrelated to the outcome once conditioning on all other controls. Standard errors are clustered at the debate level.
@@
In our empirical analysis, we take a different approach. Rather than controlling for a large set of variables as in (2), we restrict the sample to observations within a narrow time window of $[-3,+3]$ days relative to the debate. This localized approach allows us to focus on a period immediately surrounding the debate, where we define the pre-debate days as the control group ($D=0$) and the debate day along with the post-debate days as the treatment group ($D=1$).
We emphasize that we do not need to control for additional covariates, because, unlike the usual exogeneity assumption given covariates, the MTR and MTS assumptions do not require covariate adjustments for their validity, and they appear to be natural in this design as we discussed earlier. Specifically, MTR holds at the observational level, while MTS is likely to hold unconditionally. Since a debate provides information, “being undecided” is less likely in the presence of a debate, and hence MTR follows. Also, respondents surveyed later in the election cycle are naturally closer to the election day, which increases the likelihood of consistent voting behavior. Therefore, MTS follows as well. We remark that our approach does not rely on any functional form or exogeneity assumptions.
Table 1. Upper Bounds on the Persuasion Rates (in Percentage)
ATE APR ($\theta$) R-APR ($\theta^{(r)}$)
Est. SE UCB Est. SE UCB Est. SE UCB
All 0.68 0.56 1.60 3.37 2.76 7.91 0.84 0.69 1.98
Age $<$ 50 1.06 0.85 2.45 4.83 3.81 11.10 1.34 1.06 3.09
Age $\geq$ 50 0.40 0.58 1.36 2.13 3.09 7.22 0.49 0.71 1.66
U.S. 0.65 0.58 1.60 5.09 4.45 12.42 0.74 0.65 1.81
U.K. 0.61 0.90 2.08 3.69 5.43 12.63 0.72 1.06 2.46
Notes: The table presents the upper bound estimates (Est.), their standard errors (SE), and the 95% upper confidence bounds (UCB) for the average treatment effect (ATE), the average persuasion rate (APR) and its reverse version (R-APR), all expressed as percentages. Standard errors are clustered at the debate level, and the UCBs are computed using asymptotic normal approximations via the delta method.
@@
Table 1 summarizes our empirical findings. In the first row, labeled “All,” we consider the full sample. The first three columns show the upper bound on the average treatment effect (ATE), its standard error, and the 95% upper confidence bound (UCB). The estimated upper bound of ATE is 0.7%, with a UCB of 1.6%. These results align closely with those in LP, although we rely on a notably different estimation method. Since the share of TP respondents is identical to ATE in our setting, these small values imply that very few voters would have consistent vote choices if only if they watched TV debates.
@@
The middle three columns in Table 1 report the upper bounds on the average persuasion rate (APR, denoted by $\theta$), its standard error, and its 95% UCB, while the last three columns present analogous estimates for the reverse average persuasion rate (R-APR, $\theta^{(r)}$). Recall that APR and R-APR can be written as
$\textrm{APR}=\frac{\textrm{ATE}}{\textrm{share of TP}+\textrm{share of NP}},\quad\textrm{R-APR}=\frac{\textrm{ATE}}{\textrm{share of TP}+\textrm{share of AP}},$
where the denominators are less than one. Therefore, both APR and R-APR are larger than ATE, resulting in higher upper bounds. Here, our APR estimate exhibits a greater degree of amplification than R-APR, because the share of NP voters is smaller than that of AP voters. Economic interpretation of APR and R-APR in this application can be explained as follows. First, APR starts from those who would not have shown vote choice consistency without watching a debate, and it asks how many of them would have changed their behavior if they had watched a debate. In contrast, R-APR starts from those who would have shown vote choice consistency after watching a debate, and it asks how many of them would have been different if it were not for the debate. The former appears to be small, but the latter looks even smaller. Finally, we remark that the 95% confidence intervals are all one-sided because the lower bounds are trivially zero, and only the upper bounds need to be estimated (see theorem 1).
Subsequent rows of table 1 break down the analysis by subpopulations—respondents under 50, those 50 and older, and samples from the United States and the United Kingdom. Younger voters generally exhibit higher upper bounds than older voters, and the U.S. sample shows the largest upper bound on APR, suggesting a smaller share of APs in that subpopulation.
@@
Overall, our findings indicate that TV debates exert, at best, a modest influence on vote choice consistency, echoing LP’s results. However, when focusing on APR (which excludes AP voters from the relevant population), the potential scope for debate effects is larger. Even so, the evidence remains suggestive because (i) the identified parameter is merely an upper bound on the true persuasive effect, and (ii) the standard errors associated with APR are relatively large.
@@ 6. Conclusions
We first articulate the empirical contents of the MTR and MTS restrictions, after which we discuss identification of $\mathbb{P}\{\mathrm{NP}\}$, $\mathbb{P}\{\mathrm{TP}\}$, $\mathbb{P}\{\mathrm{AP}\}$, $\theta$, and $\theta^{(r)}$.
@@
Here, assumption B has already been imposed. Assumption D can be formulated by
@@
where the denominators are not equal to zero by assumption A. Assumptions B and D do not impose any other restrictions on the $q_{j}$’s.
For $(y,d)\in\{0,1\}^{2}$, let $P_{yd}:=\mathbb{P}(Y=y,D=d)$, which are directly identified from the data. Then,
$\begin{split}q_{1}+q_{2}&=P_{00},\\ q_{4}&=P_{01},\end{split}\qquad\qquad\begin{split}q_{3}&=P_{10},\\ q_{5}+q_{6}&=1-P_{00}-P_{10}-P_{01},\end{split}$
where $\sum_{j=1}^{6}q_{j}=1$. Therefore, only two of the $q_{j}$’s are undetermined. Specifically, since $q_{6}=1-\sum_{j=1}^{5}q_{j}$, the last equation is redundant, and we can write
$q_{1}=P_{00}-q_{2},\quad q_{3}=P_{10},\quad q_{4}=P_{01},\quad q_{6}=P_{11}-q_{5}.$ (5)
That is, we have expressed all the $q_{j}$’s in terms of $q_{2}$ and $q_{5}$, where $0\leq q_{2}\leq P_{00}$ and $0\leq q_{5}\leq P_{11}$. Using equation 5, equations 3 and 4 can be rewritten as
@@
$\frac{P_{11}}{P_{01}+P_{11}}$ $\geq\frac{q_{2}+P_{10}}{P_{00}+P_{10}},$ (7)
from which we obtain
@@
write $\mathbb{P}\{\mathrm{NP}\}$, $\mathbb{P}\{\mathrm{TP}\}$, and $\mathbb{P}\{\mathrm{AP}\}$ in terms of $q_{2}$ and $q_{5}$ as before, so that
$\mathbb{P}\{\mathrm{NP}\}=P_{00}+P_{01}-q_{2},$$\;\mathbb{P}\{\mathrm{TP}\}=q_{2}+q_{5},\;\mathbb{P}\{\mathrm{AP}\}=P_{10}+P_{11}-q_{5},$
which is from equation 5. Recall that the bounds on $q_{2}$ and $q_{5}$ are a rectangle (see equations 8 and 9). Thus, the sharp bounds on $\mathbb{P}\{\mathrm{NP}\}$, $\mathbb{P}\{\mathrm{TP}\}$, and $\mathbb{P}\{\mathrm{AP}\}$ are
$\mathbb{P}\{\mathrm{NP}\}$ $\in\left[\mathbb{P}(Y=0\mid D=1),\mathbb{P}(Y=0)\right],$
$\mathbb{P}\{\mathrm{AP}\}$ $\in\left[\mathbb{P}(Y=1\mid D=0),\mathbb{P}(Y=1)\right],$
while $\mathbb{P}\{\mathrm{TP}\}=1-\mathbb{P}\{\mathrm{NP}\}-\mathbb{P}\{\mathrm{AP}\}$. ∎
@@
Therefore, we can obtain the upper bounds on $\theta$ and $\theta^{(r)}$ by maximizing equations 10 and 11, respectively, subject to equations 8 and 9; the lower bounds are similar.
First, consider $\theta$. The sharp lower bound on $\theta$ can be obtained by
[…]
Since $Q_{\theta}(\cdot,q_{5})$ is monotonic, it suffices to solve
[…]
from which it follows that the lower bound on $\theta$ is $Q_{\theta}(0,0)=0$. Similarly, for the upper bound on $\theta$, we consider
@@
$Q_{\theta}\left(P_{00}-\frac{P_{01}(P_{10}+P_{00})}{P_{01}+P_{11}},\ P_{11}-\frac{P_{10}(P_{01}+P_{11})}{P_{00}+P_{10}}\right).$
Let $Q_{d}:=P_{0d}+P_{1d}$, and simplifying this leads to
$\frac{P_{00}-\frac{P_{01}Q_{0}}{Q_{1}}+P_{11}-\frac{P_{10}Q_{1}}{Q_{0}}}{1-P_{10}-\frac{P_{10}Q_{1}}{Q_{0}}}$ $=\frac{Q_{0}P_{00}-\frac{P_{01}Q_{0}^{2}}{Q_{1}}+Q_{0}P_{11}-P_{10}Q_{1}}{Q_{0}-P_{10}}$
$=\frac{Q_{0}P_{00}-\frac{P_{01}Q_{0}^{2}}{Q_{1}}+Q_{0}P_{11}+P_{10}Q_{0}-P_{10}}{Q_{0}-P_{10}}$
$=\frac{\frac{1}{Q_{1}}\left(Q_{0}Q_{1}P_{00}-P_{01}Q_{0}^{2}+Q_{0}Q_{1}P_{11}+Q_{0}Q_{1}P_{10}\right)-P_{10}}{Q_{0}-P_{10}}$
$=\frac{\frac{1}{Q_{1}}\left\{Q_{0}Q_{1}(1-P_{01})-P_{01}Q_{0}^{2}\right\}-P_{10}}{Q_{0}-P_{10}}$
$=\frac{\frac{Q_{0}}{Q_{1}}(Q_{1}-P_{01})-P_{10}}{Q_{0}-P_{10}}$
$=\frac{\frac{Q_{0}}{Q_{1}}P_{11}-P_{10}}{Q_{0}-P_{10}}$
$=\frac{P_{11}/Q_{1}-P_{10}/Q_{0}}{1-P_{10}/Q_{0}}.$

### v3

@@ 5. Bounding the Probabilities of Causation
$\mathrm{PS}$ $=\mathbb{P}\{Y(1)=1\mid Y=0,D=0\},$
$\mathrm{PN}$ $=\mathbb{P}\{Y(0)=0\mid Y=1,D=1\},$
$\mathrm{PNS}$ $=\mathbb{P}\{Y(1)=1,Y(0)=0\},$
@@
$\mathrm{PS}$ $=\frac{\mathbb{P}\{\mathrm{TP},D=0\}}{\mathbb{P}\{\mathrm{TP},D=0\}+\mathbb{P}\{\mathrm{NP},D=0\}},$
$\mathrm{PN}$ $=\frac{\mathbb{P}\{\mathrm{TP},D=1\}}{\mathbb{P}\{\mathrm{TP},D=1\}+\mathbb{P}\{\mathrm{AP},D=1\}},$
$\mathrm{PNS}$ $=\mathbb{P}\{\mathrm{TP}\}.$
Under Assumptions˜A, C and B, we have that $\theta=\mathrm{PS}$, $\theta^{(r)}=\mathrm{PN}$, and $\mathrm{PNS}=\mathbb{P}(Y=1\mid D=1)-\mathbb{P}(Y=1\mid D=0)$, as initially noted by Pearl (1999). More interestingly, under Assumptions˜A, B and D, we have the following results.
@@
Suppose that Assumptions˜A, B and D hold. Then, the sharp identified intervals on $\mathrm{PS}$, $\mathrm{PN}$, and $\mathrm{PNS}$, respectively, are given by
$0\leq\mathrm{PS}\leq\theta_{U},\;0\leq\mathrm{PN}\leq\theta^{(r)}_{U},\;\text{ and }\;0\leq\mathrm{PNS}\leq\mathbb{P}(Y=1\mid D=1)-\mathbb{P}(Y=1\mid D=0),$
@@
The sharp bounds on PNS are immediate from Theorem˜1 because $\mathbb{P}\{\mathrm{TP}\}$ equals ATE under MTR. The main point here is that the sharp identified bounds on $\mathrm{PS}$ and $\mathrm{PN}$ under Assumptions˜A, B and D coincide with those on $\theta$ and $\theta^{(r)}$, respectively.
6. Bounds under Alternative Assumptions
In this section, we derive the identified regions for the persuasion rate $\theta$ and the reverse persuasion rate $\theta^{(r)}$ under relaxed assumptions. We consider three scenarios: (i) no monotonicity assumptions; (ii) MTR only; and (iii) MTS only.
Theorem 3.
Let Assumption˜A hold. The sharp identified intervals for $\theta$ and $\theta^{(r)}$, respectively, are given as follows:
  1. (i)
    With no other assumptions: $\theta\in[0,1]$ and $\theta^{(r)}\in[0,1]$.
  2. (ii)
    With Assumption˜B only: $\theta\in[0,1]$ and $\theta^{(r)}\in[0,1]$.
  3. (iii)
    With Assumption˜D only: $\theta\in[0,\overline{\theta}_{\text{MTS}}]$ and $\theta^{(r)}\in[0,\overline{\theta}^{(r)}_{\text{MTS}}$], where
    $\overline{\theta}_{\text{MTS}}$ $:=\min\left\{\frac{\mathbb{P}(Y=1\mid D=1)}{\mathbb{P}(Y=1,D=1)+\mathbb{P}(Y=0,D=0)},\ 1\right\},$
    $\overline{\theta}^{(r)}_{\text{MTS}}$ $:=\min\left\{\frac{\mathbb{P}(Y=0\mid D=0)}{\mathbb{P}(Y=1,D=1)+\mathbb{P}(Y=0,D=0)},\ 1\right\}.$
Proof.
See the online appendix. The argument closely parallels the proof of Theorem˜1. ∎
Theorem˜3 shows that the MTR assumption alone provides no identifying power for either $\theta$ or $\theta^{(r)}$, whereas MTS alone may yield non-trivial bounds. Intuitively, the role of MTR is restricted to relating the joint probabilities of the potential outcomes to their marginal distributions, whereas MTS has identifying power for the marginals of the potential outcomes themselves.
Furthermore, note that
$\overline{\theta}_{\text{MTS}}<1$ $\Longleftrightarrow\mathbb{P}(Y=1,D=1)<\mathbb{P}(D=1)\{\mathbb{P}(Y=1,D=1)+\mathbb{P}(Y=0,D=0)\}$
$\Longleftrightarrow\mathbb{P}(Y=0,D=0)>\mathbb{P}(D=0)\{\mathbb{P}(Y=1,D=1)+\mathbb{P}(Y=0,D=0)\}$
$\Longleftrightarrow\overline{\theta}^{(r)}_{\text{MTS}}=1.$
Thus, whenever MTS yields a non-trivial upper bound on $\theta$, the sharp bound on $\theta^{(r)}$ becomes trivial. In contrast, we have
$\overline{\theta}_{\text{MTS}}\geq\theta_{U}\quad\text{and}\quad\overline{\theta}^{(r)}_{\text{MTS}}\geq\theta^{(r)}_{U}$
by construction, where both $\theta_{U}$ and $\theta^{(r)}_{U}$ can be non-trivial. We therefore adopt the joint assumption of MTR and MTS as our preferred identifying restrictions.
We now turn to the sharp identified regions for $\mathrm{PS}$, $\mathrm{PN}$, and $\mathrm{PNS}$ under the three scenarios. Define
$\overline{\mathrm{PNS}}_{\text{NA}}$ $:=\mathbb{P}(Y=1,D=1)+\mathbb{P}(Y=0,D=0),$
$\overline{\mathrm{PNS}}_{\text{MTS}}$ $:=\min\{\mathbb{P}(Y=0\mid D=0),\ \mathbb{P}(Y=1\mid D=1)\},$
$\overline{\mathrm{PS}}_{\text{MTS}}$ $:=\min\left\{\frac{\mathbb{P}(Y=1\mid D=1)}{\mathbb{P}(Y=0\mid D=0)},\ 1\right\},$
$\overline{\mathrm{PN}}_{\text{MTS}}$ $:=\min\left\{\frac{\mathbb{P}(Y=0\mid D=0)}{\mathbb{P}(Y=1\mid D=1)},\ 1\right\}.$
Theorem 4.
Let Assumption˜A hold. The sharp identified intervals for $\mathrm{PS}$, $\mathrm{PN}$, and $\mathrm{PNS}$, respectively, are given as follows:
  1. (i)
    With no other assumptions: $\mathrm{PS}\in[0,1]$, $\mathrm{PN}\in[0,1]$, and $\mathrm{PNS}\in[0,\overline{\mathrm{PNS}}_{\text{NA}}]$.
  2. (ii)
    With Assumption˜B only: $\mathrm{PS}\in[0,1]$, $\mathrm{PN}\in[0,1]$, and $\mathrm{PNS}\in[0,\overline{\mathrm{PNS}}_{\text{NA}}]$.
  3. (iii)
    With Assumption˜D only: $\mathrm{PS}\in[0,\overline{\mathrm{PS}}_{\text{MTS}}]$, $\mathrm{PN}\in[0,\overline{\mathrm{PN}}_{\text{MTS}}]$, and $\mathrm{PNS}\in[0,\overline{\mathrm{PNS}}_{\text{MTS}}]$.
Proof.
Case (i) is given in Tian and Pearl (2000, Section 4.2.1), and case (ii) is given in Tian and Pearl (2000, Section 4.4.1). For case (iii), see the online appendix. The argument closely parallels the proof of Theorem˜2. ∎
As in the case of $\theta$ and $\theta^{(r)}$, the MTR assumption alone does not have any identification power for any of the probabilities of causation. Our results under MTS only, i.e., Theorem˜4(iii), relate to the bounds derived by Tian and Pearl (2000, Theorem 1) under exogeneity. Comparing the two sets of the results, we find that the upper bounds yielded by MTS alone are identical to those under exogeneity. The additional identification power provided by exogeneity applies exclusively to the lower bounds.
Finally, Table˜1 summarizes the hierarchy of bounds for the persuasion rates and the probabilities of causation under different combinations of assumptions.
@@ 7. Empirical Example
In this section, we revisit the findings of LP to reevaluate the effectiveness of TV debates in persuading (undecided) voters. The debates examined in LP took place between 5 and 44 days (on average 24 days) before the election, precisely when vote choice consistency tends to increase most rapidly. Utilizing an event-study design based on 331,000 respondent-debate-election observations, LP isolate the impact of debates occurring one to three days post-event relative to the broadcast day. Their specification controls for secular time trends, sociodemographic covariates, and debate-by-election fixed effects, assuming the specific timing of the debate is exogenous conditional on these factors. See Appendix˜S-1 for the full econometric specification. Interestingly, LP find no significant differences in how voters’ choices evolved before versus after debate periods, rejecting any effect larger than $0.5$ percentage points. This null result holds across diverse voter types, different election contexts, and even tight, uncertain races.
@@
In our empirical analysis, we adopt an alternative strategy. Rather than controlling for the extensive set of variables used in LP, we restrict the sample to a narrow window of [3,+3][-3,+3] days relative to the debate. This localized approach focuses on the period immediately surrounding the event. We define the pre-debate days as the control group ($D=0$) and the debate day together with the post-debate days as the treatment group ($D=1$). Panel B of Figure˜1 provides a visualization of the treatment definition ($D$) and the sample window.
Crucially, our method does not require conditioning on additional covariates. Unlike standard exogeneity assumptions, the MTR and MTS assumptions in this setting do not rely on covariate adjustment for their validity. MTR is imposed at the individual level, while MTS will be motivated without controlling for covariates. Therefore, our approach avoids reliance on functional form restrictions or strict exogeneity assumptions.
First, consider the plausibility of MTR, for which it is helpful to recall that the outcome variable is vote choice consistency defined by the agreement between respondents’ reported voting intentions prior to the election and their realized vote choices reported after the election; if the voter is undecided before the election, then the outcome is coded as inconsistent voting whatever the final vote is. Panel A of Figure˜1 explains how the potential outcomes would be coded in several scenarios. The idea is that because debates provide information, the likelihood of a voter remaining undecided should decrease in the presence of a debate. For example, imagine a voter who would vote for Candidate A before a debate, would become disappointed and undecided after watching the debate, and would end up just not voting at all as a result. If this voter is observed before the debate, then $Y(0)=0$ will be observed, whereas if she is surveyed after the debate, then $Y(1)=0$ will be observed. Therefore, MTR is still satisfied in this case. From this perspective, information revealed during a debate (including details about candidates’ traits such as health or competence) helps voters refine and finalize their preferences prior to voting. Even if such information leads some voters to revise their initial intentions, this does not violate MTR, provided that debate exposure does not induce voters to become “confused” about their own preferences.
Next, regarding the plausibility of MTS, respondents surveyed later in the electoral cycle are closer to the election day and thus more likely to exhibit consistent voting behavior on average. This supports the unconditional MTS assumption. Further, our use of a $\pm$3-day estimation window around the debate adds extra support for its plausibility. While respondent composition might shift over a long campaign (e.g., urban voters being surveyed earlier than rural voters), such structural shifts are implausible within such a short period. We verify this by showing that observable covariates are balanced between the pre- and post-debate groups. Specifically, when we regress the treatment indicator on the sociodemographic controls used in the paper (gender, age, income quartiles, employment status, and education), the resulting p-value for the joint F-test of significance (with standard errors clustered at the debate level) is 0.445. This failure to reject the null hypothesis supports the validity of the unconditional MTS assumption in our local setting.
@@
Table˜2 presents the estimation results divided into two panels. Panel A reports the estimates and upper confidence bounds of (the sharp upper bounds on) the Average Treatment Effect (ATE), the Average Persuasion Rate (APR, denoted as $\theta$), and the Reverse Average Persuasion Rate (R-APR, denoted as $\theta^{(r)}$). In the first row, labeled “All,” we consider the full sample ($N=42,136$) and find that the estimated impact of TV debates is small. The point estimate of the upper bound on ATE is 0.68 percentage points (standard error 0.56), with a 95% upper confidence bound (UCB) of 1.60%. These results align closely with those in LP, although we rely on a notably different methodology. Since the share of Treatment-Persuadable (TP) respondents is identical to ATE in our setting, these small values imply that very few voters’ consistency relies exclusively on watching TV debates.
@@
The middle three columns in Panel A report the upper bounds on APR, its standard error (SE), and its 95% UCB, while the last three columns present analogous estimates for R-APR. Recall that these rates rescale the share of TP relative to different denominators:
[…]
Because both denominators are less than one, APR and R-APR are mechanically no less than ATE. Specifically, APR is estimated at 3.37% with a UCB of 7.91%, while R-APR is estimated at 0.84% with a UCB of 1.98%. The APR estimate exhibits greater amplification than that of R-APR because the share of Never-Persuadable (NP) voters is significantly smaller than that of Already-Persuaded (AP) voters. This will be confirmed in Panel B of Table˜2.
Economically, these measures answer different counterfactual questions. APR considers those who would not have been consistent without the debate and asks what fraction would change their behavior if they watched it. In contrast, R-APR considers those who were consistent after watching the debate and asks what fraction would have differed without it. The former effect is small, but the latter is even smaller. We note that the reported 95% confidence intervals are one-sided because the lower bounds are trivially zero, so only the upper bounds require estimation (see Theorem˜1).
Subsequent rows of Panel A of Table˜2 reveal notable heterogeneity across demographic groups and countries. Younger voters (Age <50<50) appear more susceptible to persuasion than older voters: the estimated sharp upper bound on ATE for the under-50 group is 1.06% (UCB 2.45%), compared to just 0.40% (UCB 1.36%) for the over-50 group. Similarly, APR for younger viewers is more than double that of older viewers (4.83% versus 2.13%). When disaggregated by country, the results for the U.S. and the U.K. are qualitatively similar, though the U.S. sample shows a slightly higher persuasion rate: APR in the U.S. is estimated at 5.09% (UCB 12.42%), compared to 3.69% (UCB 12.63%) in the U.K. However, the standard errors for these subsamples are larger, indicating less precision than in the pooled analysis.
Panel B of Table˜2 provides insight into the shares of different types: the vast majority of the population belongs to the AP or NP categories. For the full sample, the estimated proportion of AP types is bounded between 79.88% and 80.25%, while the proportion of NP types lies between 19.44% and 19.75%. The tight identification region confirms that approximately 80% of the viewing population was already committed to the outcome prior to the debate. This “ceiling effect” is more pronounced in the U.S. subsample, where the lower bound on the share of the AP type is 87.30%, slightly higher than the U.K.’s 83.59%.
@@
Overall, our findings indicate that TV debates exert, at best, a modest influence on vote choice consistency, echoing LP’s results. However, when focusing on APR—which excludes AP voters from the relevant population—the potential scope for debate effects is larger. Even so, this evidence remains suggestive because (i) the identified parameter is merely an upper bound on the true persuasive effect, and (ii) the standard errors associated with the APR estimates are relatively large.
@@ 8. Conclusions
Finally, we summarize a practical roadmap for empirical implementation, as illustrated in Figure 2. We consider a baseline setting with binary outcomes, binary treatments, and observed covariates under the MTR assumption. The first step is to assess data completeness. When outcomes or treatments are missing for some units, the sample selection bounds of Possebom and Riva (2025) are applicable under treatment exogeneity. In the absence of sample selection, the appropriate estimator depends on the treatment assignment mechanism. Under exogenous treatment, persuasion rates are point identified by sample analogs, as shown in Lemma 1. When treatment is endogenous, identification requires additional structure, such as instrumental variables (Jun and Lee, 2023; Yu, 2025) or panel data combined with parallel trends assumptions (Jun and Lee, 2024b). If neither valid instruments nor panel data are available, the bounds developed in this paper remain informative under the extra assumption of MTS.
An important direction for future research is to broaden the current framework to settings beyond the reach of existing methods, including cases featuring both treatment endogeneity and sample selection, and to develop identification and inference procedures that remain informative under such combined challenges.
@@
We first articulate the empirical contents of MTR and MTS, after which we discuss identification of $\mathbb{P}\{\mathrm{NP}\}$, $\mathbb{P}\{\mathrm{TP}\}$, $\mathbb{P}\{\mathrm{AP}\}$, $\theta$, and $\theta^{(r)}$.
@@
Here, Assumption˜B has already been imposed. Assumption˜D can be formulated by
@@
where the denominators are not equal to zero by Assumption˜A. Assumptions˜B and D do not impose any other restrictions on the $q_{j}$’s.
For $(y,d)\in\{0,1\}^{2}$, let $P_{yd}:=\mathbb{P}(Y=y,D=d)\in(0,1)$, which are directly identified from the data. Then,
$\begin{split}q_{1}+q_{2}&=P_{00},\\ q_{4}&=P_{01},\end{split}\qquad\qquad\begin{split}q_{3}&=P_{10},\\ q_{5}+q_{6}&=1-P_{00}-P_{10}-P_{01}.\end{split}$
Therefore, only two of the $q_{j}$’s are undetermined: i.e., using $q_{6}=1-\sum_{j=1}^{5}q_{j}$, we can write
$q_{1}=P_{00}-q_{2},\quad q_{3}=P_{10},\quad q_{4}=P_{01},\quad q_{6}=P_{11}-q_{5},$ (4)
where $0\leq q_{2}\leq P_{00}$ and $0\leq q_{5}\leq P_{11}$. Using Equation˜4, Equations˜2 and 3 can be rewritten as
@@
$\frac{P_{11}}{P_{01}+P_{11}}$ $\geq\frac{q_{2}+P_{10}}{P_{00}+P_{10}}.$ (6)
Therefore, the restrictions we have from Assumptions˜A, B and D are given by Equation˜4 with
@@
we have
$\mathbb{P}\{\mathrm{NP}\}=P_{00}+P_{01}-q_{2},$ $\;\mathbb{P}\{\mathrm{TP}\}=q_{2}+q_{5},\;\mathbb{P}\{\mathrm{AP}\}=P_{10}+P_{11}-q_{5}.$
Therefore, the sharp upper (lower) bound on $\mathbb{P}\{\mathrm{TP}\}$ can be obtained by maximizing (minimizing) $q_{2}+q_{5}$ subject to the constraints in (7) and (8). The other cases are similar. ∎
@@
Therefore, we can obtain the upper bounds on $\theta$ and $\theta^{(r)}$ by maximizing Equations˜9 and 10, respectively, subject to the constraints in (7) and (8); the lower bounds are similar. Below we focus on the upper bound on $\theta$.
Consider
@@
$Q_{\theta}\left(P_{00}-\frac{P_{01}(P_{10}+P_{00})}{P_{01}+P_{11}},\ P_{11}-\frac{P_{10}(P_{01}+P_{11})}{P_{00}+P_{10}}\right).$ (11)